# Superconductivity in Magnetars: Exploring Type-I and Type-II States in Toroidal Magnetic Fields

### v1

@@ I Introduction
The onset of superconducting and superfluid states in compact stars, particularly neutron stars (NSs), occurs shortly after their formation, as they cool below the critical temperature $T_{c}\sim 10^{9}\mathrm{~{}K}$, typically within hours of their birth in a supernova explosion Page et al. (2006, 2011); Blaschke et al. (2013). Finite-temperature effects, near but below the critical temperature, are significant, particularly in the context of neutrino emission from compact stars. Moreover, the proton fluid in the core of a compact star is highly degenerate, with a Fermi temperature $T_{F}\geq 10^{11}\mathrm{~{}K}\gg T_{c}$ in the relevant density range. This strong degeneracy condition is essential for the applicability of the microscopic theory of superconductivity Abrikosov (1988). It implies that the superconductor is well within the BCS regime of pairing with a coherence length much larger than the interparticle distance Sedrakian and Clark (2019).
@@ II Critical temperature and critical magnetic fields of superconductors
Fig. 1 shows the proton gap $\Delta$ as a function of density for several temperature values, calculated from Eq. (1). It is seen that the gap, and, therefore, the superconducting area within the star, grows as it cools from $T_{c}\sim 10^{10}$ K down to $T\sim 10^{8}$ K during the neutrino cooling era which takes about $10^{4}$ years Page et al. (2006). Note that the protons are bound in clusters inside the crust of a magnetar, therefore, mesoscopic and bulk superconducting effects are not present for densities $n\leq 0.5n_{s}$ corresponding to the crust-core boundary.
Superconductors on mesoscopic scales are characterized by two scales - the penetration depth $\lambda$ of the magnetic field in the superconductor and the coherence length $\xi$ characterizing the size of a Cooper pair. Their electrodynamics can be treated phenomenologically in terms of the GL theory, which contains as a key input the GL parameter $\kappa=\lambda/\xi$ Abrikosov (1988). In the following, we use $H$ for local magnetic field intensity ($H$-field) and $B$ for local magnetic field induction, which includes the response of the superconductor to an applied field, see Refs. Sedrakian and Sedrakian (1995); Haber and Schmitt (2017). For $\kappa>1/\sqrt{2}$, the surface energy between the normal and superconducting states is negative and for the local field induction between the lower and upper critical fields: $H_{c1}\leq H\leq H_{c2}$, superconductor sustains quantized fluxtubes which carry a quantum of circulation. For $H>H_{c2}$ superconductivity breaks down. Conversely, for $H<H_{c1}$, the field is supposed to be expelled from the superconducting regions; however, metastable fluxtube featuring state, in this case, was conjectured in Ref. Baym et al. (1969), as the flux expulsion timescale is rendered very long by the high electric conductivity of core material. If the magnetic field is assumed to pre-exist during the onset of superconductivity, it is treated as the local field intensity $H$ to which the critical fields $H_{c1}$ and $H_{c2}$ are compared. As the density increases $\kappa$ decreases below the limiting value $1/\sqrt{2}$ in which case type-I superconducting state is preferable. In this state, normal and superconducting states form layers of normal-superconducting domains whose size and structure depend on various factors, including nucleation dynamics, magnetic history, etc.
@@ III Magnetar models with type-I and type-II superconducting regions
    • $H>H_{c2}$, normal state is preferred;
    • $H_{c1}\leq H\leq H_{c2}$, superconductor features fluxtube arrays with magnetic field induction $B\simeq H$ except when $H\to H_{c1}$.
    • $H<H_{c1}$, the Meissner state is preferred over the mixed state with $B=0$, but an (inhomogenous) lattice of fluxtubes may arise if the expulsion time is too long with $B\simeq H$.
@@
    • $H>H_{\rm cm}$, normal state is preferred;
    • $H<H_{\rm cm}$, the Meissner state with $B=0$ is preferred which may, under specific conditions, break down into domains of normal and superconducting states with $B\neq 0$.

### v2

@@ I Introduction
The onset of superconducting and superfluid states in compact stars, particularly neutron stars (NSs), occurs shortly after their formation, as they cool below the critical temperature $T_{c}\sim 10^{9}\mathrm{~{}K}$, typically within hours of their birth in a supernova explosion Page et al. (2006, 2011); Blaschke et al. (2013). Finite-temperature effects, near but below the critical temperature, are significant, particularly in the context of neutrino emission from compact stars. Moreover, the proton fluid in the core of a compact star is highly degenerate, with a Fermi temperature $T_{F}\geq 10^{11}\mathrm{~{}K}\gg T_{c}$ in the relevant density range. This strong degeneracy condition is essential for the applicability of the microscopic theory of superconductivity Abrikosov (1988). It implies that the superconductor is well within the weak-coupling regime of pairing with a coherence length much larger than the interparticle distance Sedrakian and Clark (2019).
@@ II Critical temperature and critical magnetic fields of superconductors
Fig. 1 shows the proton gap $\Delta$ as a function of density for several temperature values, calculated from Eq. (1). It is seen that the gap, and, therefore, the superconducting area within the star, grows as it cools from $T\sim 10^{10}$ K down to $T\sim 10^{8}$ K during the neutrino cooling era which takes about $10^{4}$ years Page et al. (2006). Note that the protons are bound in clusters inside the crust of a magnetar, therefore, mesoscopic and bulk superconducting effects are not present for densities $n\leq 0.5n_{s}$ corresponding to the crust-core boundary.
Superconductors on mesoscopic scales are characterized by two scales - the penetration depth $\lambda$ of the magnetic field in the superconductor and the coherence length $\xi$ characterizing the size of a Cooper pair. Their electrodynamics can be treated phenomenologically in terms of the GL theory, which contains as a key input the GL parameter $\kappa=\lambda/\xi$ Abrikosov (1988). In the following, we use $H$ for local magnetic field intensity ($H$-field) and $B$ for local magnetic field induction, which includes the response of the superconductor to an applied field, see Refs. Sedrakian and Sedrakian (1995); Haber and Schmitt (2017). For $\kappa>1/\sqrt{2}$, the surface energy between the normal and superconducting states is negative and for the local field induction between the lower and upper critical fields: $H_{c1}\leq H\leq H_{c2}$, superconductor sustains quantized fluxtubes which carry a quantum of circulation. For $H>H_{c2}$ superconductivity breaks down. Conversely, for $H<H_{c1}$, the field is supposed to be expelled from the superconducting regions; however, metastable fluxtube featuring state, in this case, was conjectured in Ref. Baym et al. (1969), as the flux expulsion timescale is rendered very long by the high electric conductivity of core material. (If the magnetic field is assumed to pre-exist during the onset of superconductivity, it is treated as the local field intensity $H$ to which the critical fields $H_{c1}$ and $H_{c2}$ are compared.) As the density increases $\kappa$ decreases below the limiting value $1/\sqrt{2}$ in which case type-I superconducting state is preferable. In this state, normal and superconducting states form layers of normal-superconducting domains whose size and structure depend on various factors, including nucleation dynamics, magnetic history, etc.
@@ III Magnetar models with type-I and type-II superconducting regions
    $H>H_{c2}$, normal state is preferred; $H_{c1}\leq H\leq H_{c2}$, superconductor features fluxtube arrays with magnetic field induction $B\simeq H$ except when $H\to H_{c1}$; $H<H_{c1}$, the Meissner state is preferred over the mixed state with $B=0$, but an (inhomogenous) lattice of fluxtubes may arise if the expulsion time is too long with $B\simeq H$.
@@
    $H>H_{\rm cm}$, normal state is preferred; $H<H_{\rm cm}$, the Meissner state with $B=0$ is preferred which may, under specific conditions, break down into domains of normal and superconducting states with $B\neq 0$.